OpenStudy (anonymous):

In your own words, explain how Hubble's Law provides evidence for the Big Bang Theory. I know what the Law is i just need help explaining it.

OpenStudy (frostbite):

One key word: The hubble time.

OpenStudy (frostbite):

I know you know the law, yet lets take a very very close look at some of its properties: Hobbles law states that there are a linear relationship between redshift and distance, coupled with a supposed linear relation between recessional velocity and redshift, mathematical we describe this as: v = H * d Now lets take a look at the equation and rewrite it a bit: v = H * d <-> H = v / d Now when we look at speed we have another equation: t = d / v (here is t time, d lenght, v velocity) We can see that the two equations look alot alike, so we need say that t =1/H = H^-1 The argument for this is the following: v = H * d d = v / H d/v = (v/H) / (H*d) = ((v/H)/d) * (1/H) = (d/d) * (1/H) = 1* (1/H)=1/H = t t=1/H=H^-1 So we now have a chance to calculate a time from the hobble law if we assum that galaxies are moving at a constant speed (not likely (constant speed = (acceleration = 0)) So we try do some calculations: H=(71km/s)/Mpc=(71km/s)/3.09*10^19km=2.30*10^18s^-1 1/H=1/2.30*10^18s^-1=4.35*10^17s=13.8*10^9 years This means that the universe is about 14 billion years according to Hubble constant.

OpenStudy (frostbite):

I don't know if this is your own words, but this is how I would discribe it.